write an equation in point slope with the given information: through (4,44) and parallel to y axis
Answer by tarungehlot(9)

write an equation in point slope with the given information: through (4,44) and parallel to the y-axis?
solution =point slope form is
y -y1 =m (x -x1)
(x1,y1)= (4,44)
so
y- 44 = m( x-4)
now this line is parallel to y axis or line y=0
parallel lines have same slopes
so
required line is
x=4